OpenStudy (anonymous):

solve 2x +7y=0 show work

OpenStudy (anonymous):

solve for ?

OpenStudy (anonymous):

csc theta ?

OpenStudy (anonymous):

the whole problem is line: 2x+7y=0 on quadrant 2 evaluate: csc theta

OpenStudy (anonymous):

the whole problem is line: 2x+7y=0 on quadrant 2 evaluate: csc theta

OpenStudy (anonymous):

but the answer is sqrt 53/2

OpenStudy (anonymous):

the whole problem is line: 2x+7y=0 on quadrant 2 evaluate: csc theta

OpenStudy (anonymous):

wat happened to the drawing? D:

OpenStudy (anonymous):

2x +7y=0 7y=-2x y= -2/7 x csc is hypontenuse/opposite

OpenStudy (anonymous):

\[\sqrt{53}/2\]

OpenStudy (anonymous):